OpenStudy (anonymous):

By what number must you multiply each side of the equation 2/3x=10 to produce the equivalent equation x=15 ?

OpenStudy (anonymous):

a 3 b 3/2 c 1/3 d -2/3

OpenStudy (anonymous):

2/3x = 10 multiply both sides by 3/2 2/3x(3/2) = 10(3/2) x = 30/2 x = 15

OpenStudy (the_fizicx99):

2/3x = 10 Isolate x so multiply is by 3, 2x = 30 /2 /2 x = 15

OpenStudy (anonymous):

so i think its gonna be d is it

OpenStudy (the_fizicx99):

.___________. seriously???? O_O

OpenStudy (anonymous):

what i thought it was d is it

OpenStudy (anonymous):

isn't there 2 answers for this

OpenStudy (the_fizicx99):

Ohh they're asking for a fraction, my bad. Still works though lol xD U.U it's not D

OpenStudy (the_fizicx99):

No because if it were -3/2 you would've been left with a negative

OpenStudy (the_fizicx99):

-2/3 *

OpenStudy (anonymous):

no there isnt two answers for this

OpenStudy (anonymous):

-2/3 is a fraction

OpenStudy (anonymous):

It can be 3 or 3/2 2/3x = 10 multiply by 3 2x = 30 x = 15 OR 2/3x = 10 multiply by 3/2 x = 30/2 x = 15 Two answers...is that right @tHe_FiZiCx99

OpenStudy (the_fizicx99):

Umm yes but it's the same because its a fraction,